OpenStudy (anonymous):

Solve the following system of equations. Answer with an ordered pair. 2x+7y=-1 4x-3y=-19 Help, please and thank you.

OpenStudy (campbell_st):

double the 1st equation then subtract the 2nd equation 4x + 14y = -2 - 3x - 3y = -19 -------------- 17y = 17 solve for y when you get y substitute into either equation to find x